人 。I'm having trouble with 3,4 and 5.

A block with mass m =6.8 kg is hung from a vertical spring. When the mass hangs in equilibrium, the spring stretches x = 0.23 m. While at this equilibrium position, the mass is then given an initial push downward at v = 5 m/s. The block oscillates on the spring without friction.

1)

What is the spring constant of the spring? 289.74N/m

2)What is the oscillation frequency? 1.039 Hz


3)After t = 0.31 s what is the speed of the block? ______m/s


4)What is the magnitude of the maximum acceleration of the block? ______m/s2

5)At t = 0.31 s what is the magnitude of the net force on the block? _____N
Answer #1

Dear student,

Find this solution, and RATE IT ,If you find it is helpful .your rating is very important to me.If any incorrectness ,kindly let me know I will rectify them soon.

Thanks for asking ..at equilibrium, KXo = mg k=spring Constant = 6.889.8 - 1997 8.2 = 289.739 N/m 0.23 2) frequency (f) = zelo lectores are zat=W